OpenStudy (idealist10):

Find the general solution of (x-2)(x-1)y'-(4x-3)y=(x-2)^3.

OpenStudy (idealist10):

@myininaya

OpenStudy (idealist10):

@beccaboo333

myininaya (myininaya):

so i assumed you tried to put in the form y'+p*y=q form

OpenStudy (idealist10):

Please help me. How do I start? Divide (x-2)?

myininaya (myininaya):

divide by (x-2)(x-1)

OpenStudy (idealist10):

Okay, I'll try and tell me what's going on. Please don't leave.

myininaya (myininaya):

you may have to some partial fractions just to let you know

OpenStudy (idealist10):

But how do I integrate (4x-3)/((x-2)(x-1))?

myininaya (myininaya):

partial fractions

OpenStudy (idealist10):

So do I multiply (x-2)(x-1)? It's x^2-3x+2.

myininaya (myininaya):

no

myininaya (myininaya):

\[\frac{3-4x}{(x-2)(x-1) }=\frac{A}{x-2}+\frac{B}{x-1}\] before you integrate you write what is on the left hand side in the right hand sides's form you must find A and B to complete that transaction

OpenStudy (idealist10):

So 4x/(x-2)-3/(x-1)? And 4x(x-1)-3(x-2)? But equal to what?

myininaya (myininaya):

... Have you ever done partial fractions before?

OpenStudy (idealist10):

So isn't it 5ln(x-2)-ln(x-1)?

myininaya (myininaya):

I will show an example: \[\frac{1}{(x-5)(x-3)}\] pretend we wanted to write this as a sum of proper fractions we can attempt to do this by trying to write that fraction in this form A/(x-5)+B/(x-3) To find what A and B are we will need to combine those fractions in put in to form that is 1/[(x-5)(x-3)] form. so let's do that \[\frac{A}{x-5}+\frac{B}{x-3}=\frac{A(x-3)+B(x-5)}{(x-5)(x-3)}=\frac{(A+B)x-3A-5B}{(x-5)(x-3)}\] but the other side is equal to 1/[(x-5)(x-3)] so that means the other can have no x's so A+B would have to be 0 and the constant term will have to be equal to 1 therefore -3A-5B=1 so we have a system of linear equations to solve: A+B=0 -3A-5B=1 ----------- To solve I will try to setup for elimination Multiply first equation by 3 3A+3B=0 -3A-5B=1 -----------now add -2B=1 B=-1/2 If B=-1/2 and A+B=0, then A=1/2 So you know that \[\frac{1}{(x-5)(x-3)} \text{ can be written as } \frac{1}{2} \frac{1}{x-5}-\frac{1}{2}\frac{1}{x-3}\] and we know how to intgrate both of those terms

myininaya (myininaya):

@Idealist10 for the integrating factor there was a negative sign in front of y term so take what you have and multiply it by -1 but don't forget the integrate factor has that base e thing so for integrate factor we should have \[v=e^{-5\ln|x-2|+\ln|x-1|}\] This can be a lot prettier

myininaya (myininaya):

Now that was just one example above from partial fractions There are other things to consider when writing a fraction composed of polynomials as a sum of fractions

myininaya (myininaya):

What I'm saying is the section on partial fractions can not be explained by one example

myininaya (myininaya):

Anyways back to what you were saying...

myininaya (myininaya):

to write your integrating factor a lot prettier you will need to recall some law of exponents and also some log properties

myininaya (myininaya):

you do recall that x^(a+b) can be written as (x^a)(x^b) ?

OpenStudy (idealist10):

Yes.

myininaya (myininaya):

do you see how that could be helpful here?

OpenStudy (idealist10):

Yes, wait a second.

OpenStudy (idealist10):

So I got v=(x-1)/(x-2)^5, is this right?

myininaya (myininaya):

Very good!

OpenStudy (idealist10):

Let me work it out then.

OpenStudy (idealist10):

Y

OpenStudy (idealist10):

Yes, I got it! Thank you so much! Wow, this is a tough problem to solve!

myininaya (myininaya):

Probably mostly because of the partial fractions part?

OpenStudy (idealist10):

Yes!

myininaya (myininaya):

I guess you are in differential equations and try to recall all the things you learned in calculus 2. I do agree I had some trouble but with practice it should become easier on what you should do when you see something like that.

myininaya (myininaya):

just practice practice it will sink in eventually